Dec. 25 1999 Owl shift report.

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Sat, 25 Dec 1999 08:24:10 -0500 (EST)

Shift Summary for owl shift, 25/Dec/99

On Shift: Ruth, Valeri.

Hours of beam available at all intensities: 6.8
Hours of beam available at nominal intensity: 6.8
Hours of E832 data (special runs not included): 5.7
Hours of special runs (laser scans, muon runs...): 0.1

Accounting of lost time:

kfrends, pipeline problems. run starts/stops, magnet polarity change.

========================================================================
Runs ended this shift:
Run 15283 - 21.0 M evts - BEAM_799
Run 15285 - 0.5 M evts - BEAM_MU_4E12
Run 15286 - in progress - BEAM_799

Tapes brought to FCC this shift:

(including tapes for online split)

UP(ABC)250-252, UPS084
========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Smooth datataking with occasional problems. From time to time were getting
scaler alarms, mostly BA related.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15283 23:45 Onsplit asks for UPLA12 again. Cleaned the drive and
put in new tape.

01:25 Trigger hangs because of HCC buffer full, kfrend
fixes it.
01:26 Mounted UPF064 on ktev4.

01:50 Trigger hangs because of HCC buffer full, kfrend.

03:00 Tapes full, run stops. Do pipeline init to fix
frequent "messages from pipeline", crate 3 did
not reset, tried to treboot it - didn't work either.
Did pipe_cleanup full and init pipeline again, this
time it worked.

03:10 Mounted UPPE10 on tps2d3 (ktev3)

15285 03:30 After changing tapes and flipping magnets start
muon run. Onsplit asks for UPLA12 on ktev3 once
more. Since this is not critical we didn't page
experts and left it as is until day shift.

03:45 Switched back to 799 mode with NM4AN NEG and
NM2SR POS. While initializing new run after fresh
start got TRD alarm: TRD2 had too high anode current
(~13 uA onspill and ~8 uA offspill on both planes).
Cycling HV fixed it.

15286 03:56 Start BEAM_799 run.

04:31 DYC53 full, run kfrend, lost peemit after it and run
kfrend again.

04:50 Mounted UPM049 on ktev2.

05:17 HCC buffer full -> kfrend.

========================================================================
Follow up Notes on Previous Shift
>- see what BobT says about the CsI "heater" temperature.

nothing so far

========================================================================
Items to Follow up on Next Shift

- see what BobT says about the CsI "heater" temperature.

- check the tape drive on ktev3.